OpenStudy (anonymous):

a bag contains 4 red 2 blue 6 green and 8 white marbles. what is the probability of selecting a green marble, replacing it, then selecting a red marble

OpenStudy (anonymous):

3/10 (green) 1/5 (red)

OpenStudy (anonymous):

\[\frac{6}{20}*\frac{4}{20}=\frac{3}{50} \]

OpenStudy (anonymous):

Since you are replacing the marble, these are independent events. To find the probability of two independent events, you simply multiply the probabilities of each occurring by each other. So, first we find the probability of selecting a green marble. Since there are 6 green marbles, and 20 marbles total, that probability is \(\dfrac{6}{20}=\dfrac{3}{10}\). Then, we find the probability of selecting a red marble. There are 4 red marbles, and still 20 total, so we have \(\dfrac{4}{20}=\dfrac{1}{5}\). So, we simply multiply the two by each other, \(\dfrac{3}{10}\cdot\dfrac{1}{5}=\dfrac{3}{50}.\)